A solar collector is 2.5m long by 2.0m wide. It is held in place by a frame of uniform width around its outside edge. If the exposed collector area is 2.5m^2 , what is the width of the frame, to the nearest tenth of a centimeter?
Answer by venugopalramana(3286)

A solar collector is 2.5m long by 2.0m wide. It is held in place by a frame of uniform width around its outside edge. If the exposed collector area is 2.5m^2 , what is the width of the frame, to the nearest tenth of a centimeter?
let the width of frame = x
length of frame=x+2.5+x=2.5+2x
width of frame =x+2+x=2x+2
area of frame & reflector assembly=(2x+2.5)(2x+2)=4x^2+9x+5
area of collector=2.5*2=5
exposed area =4x^2+9x+5-5=2.5
4x^2+9x-2.5=0
8x^2+18x-5=0
x={-18+sqrt(18^2+4*5*8)}/16=1/4 metre=100/4 cm = 25 cm.